OpenStudy (anonymous):

Which graph is the line (y+1)=3/5(x-3)?

OpenStudy (anonymous):

We need to turn it into y=mx+b form so that it is easier to identify the graph can you please do that? :) Also please post the images of the graph ^_^

OpenStudy (anonymous):

@Moe95

OpenStudy (anonymous):

\[\huge~\rm~ (y+1)=3/5(x-3)\] ok turn this into slope-intercept form please: )

OpenStudy (anonymous):

so this one?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

Thanks! :)